# Constraints from CMB lensing tomography with projected bispectra

### v1

@@ 4.2.2 Probe complementarity
Figure 5: Constraints on $\sigma_{8}$ and $\Omega_{m}$ from three data combinations: $\Phi^{ggg}_{LL\ell}+C_{\ell}^{gg}$ (blue contours in top panel), $\Phi^{ggg}_{LL\ell}+C_{\ell}^{g\kappa}$ (red contours, middle panel), and $C_{\ell}^{gg}+C_{\ell}^{g\kappa}$ (yellow contours, bottom panel). The black contours show the degeneracy directions of individual power spectrum or FSB measurements, with the filled contours showing the combined constraints.. The constraints shown were obtained for redshift bin $z_{3}$.
As a pedagogical detour, before presenting a cosmological analysis of the data vector presented in the last section, let us study the ability of the galaxy bispectrum to break parameter degeneracies and improve cosmological constraints. To do so, let us use a simpler theoretical model in which we keep $\Omega_{m}$ fixed to its fiducial value (see Table 1), while varying both $\sigma_{8}$ and $b_{1}$, the two parameters that affect the overall amplitude of the clustering statistics (both $C_{\ell}$s and FSBs). For simplicity, we will also focus on the third redshift bin, $z_{3}$, alone, where the $ggg$ FSB has the highest SNR.
@@
In general, for the FSB measurements to be sufficiently complementary to $C_{\ell}^{gg}$ and $C_{\ell}^{g\kappa}$ in breaking parameter degeneracies, these scaling indices should be sufficiently different from $\alpha=4$ and $\beta=2$. Although our measurement of the $gg\kappa$ FSB is not sufficiently sensitive, we can exploit the $ggg$ FSB to break degeneracies in the $(\sigma_{8},b_{1})$ parameter plane using this rationale. Fig. 5 illustrates how, by combining $\Phi^{ggg}_{LL\ell}$ with either $C_{\ell}^{gg}$ or $C_{\ell}^{g\kappa}$, we are able to constrain $\sigma_{8}$. For the third redshift bin, the resulting constraints on $\sigma_{8}$ are [^8] We report $68\%$ confidence level intervals throughout.:
@@ 5 Discussion
We use data from four photometric DESI Luminous Red Galaxy samples, described in Zhou et al. (2023b) and White et al. (2022), and the CMB lensing convergence map obtained from Planck (Planck Collaboration et al., 2020a) to measure auto- and cross-FSBs. One of our primary results is the detection of both the galaxy bispectrum ($ggg$) at very high significance ($>30\sigma$) in all redshift bins, and the galaxy-galaxy-convergence bispectrum ($gg\kappa$) at lower significance ($>5\sigma$) in three out of the four galaxy samples used in this work. The latter being detected at a low signal-to-noise means that we are not able to use $gg\kappa$ in the second part of the paper, where we conduct a rudimentary cosmological analysis.
@@
In this first cosmological analysis including the projected bispectrum, we simply focus on three parameters – the clustering amplitude $\sigma_{8}$, the non-relativistic matter fraction $\Omega_{m}$, and the linear bias $b_{1}$ – while keeping other parameters fixed. We identify two new probe combinations that break degeneracies between $\sigma_{8}$ and $b_{1}$: combining the $ggg$ FSB measurements with any of the power spectra ($gg$ or $g\kappa$) allows us to derive complementary constraints on $\sigma_{8}$ and $\Omega_{m}$, and to test the robustness of the constraints obtained from the standard $2\times 2$-point combination against observational and modelling systematics. The new $ggg+gg$ combination is especially exciting, potentially yielding a fully local measurement of $\sigma_{8}$ without resorting to using CMB lensing data to break the $\sigma_{8}$-bias degeneracy.
@@
The addition of the $ggg$ bispectrum to both power spectra ($ggg+gg+g\kappa$ combination) leads to a modest improvement of $\sim 10\%$ on the $\sigma_{8}$ constraints. Further improvement may be achievable with more sensitive datasets, allowing us to use the $gg\kappa$ bispectrum, or pushing to scales smaller than the conservative cuts assumed here. We also find that the $g\kappa$ cross-spectrum is indispensable to obtain tight constraints on $\sigma_{8}$ (and similarly with the $gg$ auto-spectrum for $\Omega_{m}$). All combinations studied yield consistent results: $gg+g\kappa$ and $ggg+g\kappa$ are consistent between themselves at the $1\sigma$ level in $\sigma_{8}$, and at the $2\sigma$ level with $ggg+gg$. All combinations give $\sigma_{8}$ values consistently lower than inferred from Planck measurements, in agreement with previous CMB lensing tomography studies conducted on the same data (White et al., 2022, using the standard $gg+g\kappa$ combination with a different modelling approach). In particular, we find that the $ggg+gg$ combination favours even lower $\sigma_{8}$ values than other probe combinations at low redshift. We argue, however, that this result is highly dependent on some of our analysis choices, and thus refrain from fully interpreting this result in terms of a cosmological tension.
@@
Indeed, our modelling of the bispectrum comes with various caveats: in this first approach, we only model the bispectrum at the tree-level (a decision which we must compensate for by implementing conservative scale cuts) and, critically, choose to use empirical coevolution relations to express higher-order bias parameters ($b_{2}$ and $b_{s}$). These coevolution relations play an important role in our final results, as they essentially dictate the effective scaling of the bispectrum with $b_{1}$ – and thus, how efficiently it can break parameter degeneracies when combined with power spectrum measurements. Upcoming work (Verdiani et al., 2025) will focus on providing a better model for the bispectrum, in combination with a modelling of the power spectrum at the 1-loop level. This will allow us to push to smaller scales and to constrain the higher-order bias parameters directly from the data, without having to resort to using coevolution relations. In this scenario, the bispectrum is a critical tool to improve the constraints on $b_{2}$ and $b_{s}$, potentially leading to a relatively larger improvement on the final cosmological constraints obtained from power spectra than we report here.
@@ Appendix A Effect of using coevolution relations
Throughout this paper, we use coevolution relations of bias parameters to simplify our inference pipeline and redirect the constraining power of the FSB probes towards parameters of cosmological interest. For completeness, we repeat their expressions here (Lazeyras et al., 2016; Lazeyras & Schmidt, 2018):
@@
Figure 9: Plot showing constraints obtained when freeing the $b_{2}$ and $b_{s}$ parameters (light contours) as opposed to fixing them using coevolution relations (darker contours in the $b_{1}-\sigma_{8}$ block). We show the coevolution relations (Eqs. 54 and 55) as dark grey lines in the $b_{2}-b_{1}$ and $b_{s}-b_{1}$ blocks. The black dashed lines along the $b_{2}$ and $b_{s}$ rows and columns show the values of those parameters obtained via coevolution, for the published value of the linear bias for this redshift bin ($z_{2}$) – we use those as central values for the uniform priors placed on $b_{2}$ and $b_{s}$. We only show the area of parameter space allowed by the priors.
@@
We also explore the impact of allowing $b_{2}$ and $b_{s}$ to vary freely, instead of fixing them to $b_{1}$ via coevolution relations. The priors defined for these two parameters still make use of the coevolution relations, but allow a wider range of bias values: we use $\mathcal{U}[b_{2}(b_{1})-2,b_{2}(b_{1})+2]$ and $\mathcal{U}[b_{s}(b_{1})-2,b_{s}(b_{1})+2]$ as priors for $b_{2}$ and $b_{s}$ respectively, where $b_{1}$ here is the published linear bias value in the redshift bin (see Table 2), taken from White et al. (2022). Interestingly, as can be seen by comparing the light blue and light red contours in both Fig. 8 and Fig. 9, the constraints obtained in this approach are similar to those obtained when setting the non-linear biases to zero. In general, the $ggg$ FSB does not have enough constraining power on its own to fully constrain $b_{1}$, $b_{2}$ and $b_{s}$ simultaneously, as shown by the light blue, red or gray contours for the non-linear bias parameters in Fig. 9, which are largely unconstrained. In order to self-consistently constrain higher-order bias parameters from the power- and bispectrum, it will thus be crucial to push the analysis to smaller scales by modelling the power spectrum up to 1-loop and potentially moving beyond tree-level bispectra. We leave this analysis for future work.
@@ Appendix B Stochastic contributions
In this appendix we derive the expressions for the noise contributions to the $gg\kappa$ bispectrum, Eq. 34. In Appendix D of H25, we derived the equivalent term for the $ggg$ bispectrum (Eq. 33), in which case all fields measured receive a shot-noise contribution due to the discreteness of the tracer. Here, that is not the case for the third field (the CMB convergence $\kappa$). Let us see what this implies for the additional shot noise terms contributing to the $gg\kappa$ bispectrum.
@@
Now taking the expectation value over fields $g$, $\kappa$, and assuming $\langle g\rangle$ and $\langle\kappa\rangle$ to be 00 over field realisations, we obtain:
@@ Appendix C Multifield Covariance
Figure 10: Application of the multi-field covariance derivations presented in this Appendix to the galaxy and convergence data. The example data vector here contains $ggg$ FSB measurements, the $gg$ power spectrum measurement, $gg\kappa$ FSBs, and the $g\kappa$ power spectrum. We plot the correlation matrix (a normalised covariance matrix, where the diagonal is by construction unity) to better highlight off-diagonal contributions – note that we set the maximum value of the colour map to $0.5$, as essentially only the main diagonal lies above this threshold. The $N_{222}$ contributions can be seen as small brown or purple squares along the main diagonals in the four large FSB-FSB blocks (although it is barely visible in the $gg\kappa$-$gg\kappa$ blocks). The $N_{32}$ contributions show up as dark purple bands in the FSB-$C_{\ell}$ blocks. Due to the weak signal in the generalised $gg\kappa$ and $g\kappa\kappa$ FSBs, this contribution may sometimes appear as a negative contribution (light blue pixels within the $N_{32}$ bands in the $gg\kappa$-$gg$, $ggg$-$g\kappa$ and $gg\kappa$-$g\kappa$ blocks).
@@ C.1.1 $gg\kappa\times gg\kappa$ covariance
We take a look at the “$2+2+2$” decomposition since we know from the single-field case (see Appendix C.1 of H25) that it should give rise to additional contributions along the main diagonal. To keep track of individual combinations of the 6 fields involved in this covariance, we write
@@
⟨$\cancel{\langle ad\rangle\langle bc\rangle\langle ef\rangle}+\langle ad\rangle\langle be\rangle\langle cf\rangle+\langle ad\rangle\langle bf\rangle\langle ce\rangle+$
⟨$\cancel{\langle ae\rangle\langle bc\rangle\langle df\rangle}+\langle ae\rangle\langle bd\rangle\langle cf\rangle+\langle ae\rangle\langle bf\rangle\langle cd\rangle+$
⟨$\cancel{\langle af\rangle\langle bc\rangle\langle de\rangle}+\langle af\rangle\langle bd\rangle\langle ce\rangle+\langle af\rangle\langle be\rangle\langle cd\rangle$ (71)
@@ C.2.1 $gg\kappa\times g\kappa$ covariance
for simplicity, and now look at the “$3+2$” decomposition which gives rise to noticeable off-diagonal contributions in the 1-field case. This gives us the 10 following combinations:
[…]
@@ C.3 Calibrating the $f_{\mathrm{sky}}$ correction
This poses no issues in the 1-field case, where the original field and its filtered-squared version share the same footprint on the sky; the choice of rescaling factor becomes less obvious when the two fields each have their respective masks. Take Eq. 75 for instance; the expression includes $C^{gg}_{\ell}$ and $C^{g\kappa}_{\ell}$, which were measured on different effective sky areas – in this case, should the $N_{222}$ contribution be rescaled using $f_{\mathrm{sky}}=\langle w_{g}w_{g}\rangle$ or $f_{\mathrm{sky}}=\langle w_{g}w_{\kappa}\rangle$ The most conservative approach would be to always use the smallest effective sky fraction, as to not underestimate measurement errors, but this might be suboptimal.
@@
  • For the covariance of $\hat{\Phi}^{gg\kappa}_{LL\ell}$ and $\hat{\Phi}^{ggg}_{L^{\prime}L^{\prime}\ell^{\prime}}$, $N_{\tilde{\mathcal{S}}}$ is a subset of $N_{\mathcal{S}}$ by definition, since one power spectrum ($ggg$ FSB) is defined over the $g$ mask, while the other ($gg\kappa$) has for effective sky fraction the intersection of the $g$ and $\kappa$ masks. $N_{\tilde{\mathcal{S}}}$ and $N_{\mathcal{S}\cap\tilde{\mathcal{S}}}$ are therefore the same and cancel out, such that the expression must scale with $f_{\mathrm{sky}}=\langle w_{g}w_{g}\rangle$.

### v2

@@ 4.2.2 Probe complementarity
Figure 5: Constraints on $\sigma_{8}$ and $b_{1}$ from three data combinations: $\Phi^{ggg}_{LL\ell}+C_{\ell}^{gg}$ (blue contours in top panel), $\Phi^{ggg}_{LL\ell}+C_{\ell}^{g\kappa}$ (red contours, middle panel), and $C_{\ell}^{gg}+C_{\ell}^{g\kappa}$ (yellow contours, bottom panel). The black contours show the degeneracy directions of individual power spectrum or FSB measurements, with the filled contours showing the combined constraints. The constraints shown were obtained for redshift bin $z_{3}$.
As a pedagogical detour, before presenting a cosmological analysis of the data vector presented in the last section, let us study the ability of the galaxy bispectrum to break parameter degeneracies and improve cosmological constraints. To do so, let us use a simpler theoretical model in which we keep $\Omega_{m}$ fixed to its fiducial value (see Table 1), while varying both $\sigma_{8}$ and $b_{1}$, the two parameters that affect the overall amplitude of the clustering statistics (both $C_{\ell}$s and FSBs). For simplicity, we will also focus on the third redshift bin, $z_{3}$, alone, where the gggggg FSB has the highest SNR.
@@
In general, for the FSB measurements to be sufficiently complementary to $C_{\ell}^{gg}$ and $C_{\ell}^{g\kappa}$ in breaking parameter degeneracies, these scaling indices should be sufficiently different from $\alpha=4$ and $\beta=2$. Although our measurement of the $gg\kappa$ FSB is not sufficiently sensitive, we can exploit the gggggg FSB to break degeneracies in the $(\sigma_{8},b_{1})$ parameter plane using this rationale. Fig. 5 illustrates how, by combining $\Phi^{ggg}_{LL\ell}$ with either $C_{\ell}^{gg}$ or $C_{\ell}^{g\kappa}$, we are able to constrain $\sigma_{8}$. For the third redshift bin, the resulting constraints on $\sigma_{8}$ are [^9] We report $68\%$ confidence level intervals throughout.:
@@ 5 Discussion
We use data from four photometric DESI Luminous Red Galaxy samples, described in Zhou et al. (2023b) and White et al. (2022), and the CMB lensing convergence map obtained from Planck (Planck Collaboration et al., 2020a) to measure auto- and cross-FSBs. One of our primary results is the detection of both the galaxy bispectrum (gggggg) at very high significance ($30\sigma$) in all redshift bins, and the galaxy-galaxy-convergence bispectrum ($gg\kappa$) at lower significance ($5\sigma$) in three out of the four galaxy samples used in this work. The latter being detected at a low signal-to-noise means that we are not able to use $gg\kappa$ in the second part of the paper, where we conduct a rudimentary cosmological analysis.
@@
In this first cosmological analysis including the projected bispectrum, we simply focus on three parameters – the clustering amplitude $\sigma_{8}$, the non-relativistic matter fraction $\Omega_{m}$, and the linear bias $b_{1}$ – while keeping other parameters fixed. We identify two new probe combinations that break degeneracies between $\sigma_{8}$ and $b_{1}$: combining the gggggg FSB measurements with any of the power spectra ($gg$ or $g\kappa$) allows us to derive complementary constraints on $\sigma_{8}$ and $\Omega_{m}$, and to test the robustness of the constraints obtained from the standard $2\times 2$-point combination against observational and modelling systematics. The new $ggg+gg$ combination is especially exciting, potentially yielding a fully local measurement of $\sigma_{8}$ without resorting to using CMB lensing data to break the $\sigma_{8}$-bias degeneracy.
@@
The addition of the gggggg bispectrum to both power spectra ($ggg+gg+g\kappa$ combination) leads to a modest improvement of $\sim 10\%$ on the $\sigma_{8}$ constraints. Further improvement may be achievable with more sensitive datasets, allowing us to use the $gg\kappa$ bispectrum, or pushing to scales smaller than the conservative cuts assumed here. We also find that the $g\kappa$ cross-spectrum is indispensable to obtain tight constraints on $\sigma_{8}$ (and similarly with the $gg$ auto-spectrum for $\Omega_{m}$). All combinations studied yield consistent results: $gg+g\kappa$ and $ggg+g\kappa$ are consistent between themselves at the $1\sigma$ level in $\sigma_{8}$, and at the $2\sigma$ level with $ggg+gg$. All combinations give $\sigma_{8}$ values consistently lower than inferred from Planck measurements, in agreement with previous CMB lensing tomography studies conducted on the same data (White et al., 2022, using the standard $gg+g\kappa$ combination with a different modelling approach). In particular, we find that the $ggg+gg$ combination favours even lower $\sigma_{8}$ values than other probe combinations at low redshift. We argue, however, that this result is highly dependent on some of our analysis choices, and thus refrain from fully interpreting this result in terms of a cosmological tension.
@@
Indeed, our modelling of the bispectrum comes with various caveats: in this first approach, we only model the bispectrum at the tree-level (a decision which we must compensate for by implementing conservative scale cuts) and, critically, choose to use empirical coevolution relations to express higher-order bias parameters ($b_{2}$ and $b_{s}$). These coevolution relations play an important role in our final results, as they essentially dictate the effective scaling of the bispectrum with $b_{1}$ – and thus, how efficiently it can break parameter degeneracies when combined with power spectrum measurements. Upcoming work (Verdiani et al., 2025) will focus on providing a better model for the bispectrum, in combination with a modelling of the power spectrum at the 1-loop level. This will allow us to push to smaller scales and to constrain the higher-order bias parameters directly from the data, without having to resort to using coevolution relations. In this scenario, the bispectrum is a critical tool to improve the constraints on $b_{2}$ and $b_{s}$, potentially leading to a relatively larger improvement on the final cosmological constraints obtained from power spectra than we report here. Eventually, the use of hybrid approaches to model galaxy bias (e.g. Modi et al., 2020; Hadzhiyska et al., 2021) will be required to extend the range of scales over which the bispectrum may be used for cosmological inference.
@@ Appendix A Effect of using coevolution relations
Throughout this paper, we use coevolution relations of bias parameters to simplify our inference pipeline and redirect the constraining power of the FSB probes towards parameters of cosmological interest. For completeness, we repeat their expressions here (Lazeyras et al., 2016; Lazeyras and Schmidt, 2018):
@@
Figure 9: Plot showing constraints obtained when freeing the $b_{2}$ and $b_{s}$ parameters (light contours) as opposed to fixing them using coevolution relations (darker contours in the $b_{1}-\sigma_{8}$ block). We show the coevolution relations (Eqs. 54 and $5$) as dark grey lines in the $b_{2}-b_{1}$ and $b_{s}-b_{1}$ blocks. The black dashed lines along the $b_{2}$ and $b_{s}$ rows and columns show the values of those parameters obtained via coevolution, for the published value of the linear bias for this redshift bin ($z_{2}$) – we use those as central values for the uniform priors placed on $b_{2}$ and $b_{s}$. We only show the area of parameter space allowed by the priors.
@@
We also explore the impact of allowing $b_{2}$ and $b_{s}$ to vary freely, instead of fixing them to $b_{1}$ via coevolution relations. The priors defined for these two parameters still make use of the coevolution relations, but allow a wider range of bias values: we use $\mathcal{U}[b_{2}(b_{1})-2,b_{2}(b_{1})+2]$ and $\mathcal{U}[b_{s}(b_{1})-2,b_{s}(b_{1})+2]$ as priors for $b_{2}$ and $b_{s}$ respectively, where $b_{1}$ here is the published linear bias value in the redshift bin (see Table 2), taken from White et al. (2022). Interestingly, as can be seen by comparing the light blue and light red contours in both Fig. 8 and Fig. 9, the constraints obtained in this approach are similar to those obtained when setting the non-linear biases to zero. In general, the gggggg FSB does not have enough constraining power on its own to fully constrain $b_{1}$, $b_{2}$ and $b_{s}$ simultaneously, as shown by the light blue, red or gray contours for the non-linear bias parameters in Fig. 9, which are largely unconstrained. In order to self-consistently constrain higher-order bias parameters from the power- and bispectrum, it will thus be crucial to push the analysis to smaller scales by modelling the power spectrum up to 1-loop and potentially moving beyond tree-level bispectra. We leave this analysis for future work.
@@ Appendix B Stochastic contributions
In this appendix we derive the expressions for the noise contributions to the $gg\kappa$ bispectrum, Eq. 34. In Appendix D of H25, we derived the equivalent term for the gggggg bispectrum (Eq. 33), in which case all fields measured receive a shot-noise contribution due to the discreteness of the tracer. Here, that is not the case for the third field (the CMB convergence $\kappa$). Let us see what this implies for the additional shot noise terms contributing to the $gg\kappa$ bispectrum.
@@
Now taking the expectation value over fields $g$, $\kappa$, and assuming $\langle g\rangle$ and $\langle\kappa\rangle$ to be 0 over field realisations, we obtain:
@@ Appendix C Multifield Covariance
Figure 10: Application of the multi-field covariance derivations presented in this Appendix to the galaxy and convergence data. The example data vector here contains gggggg FSB measurements, the $gg$ power spectrum measurement, $gg\kappa$ FSBs, and the $g\kappa$ power spectrum. We plot the correlation matrix (a normalised covariance matrix, where the diagonal is by construction unity) to better highlight off-diagonal contributions – note that we set the maximum value of the colour map to $0.5$, as essentially only the main diagonal lies above this threshold. The $N_{222}$ contributions can be seen as small brown or purple squares along the main diagonals in the four large FSB-FSB blocks (although it is barely visible in the $g\kappa-gg$$g\kappa$ blocks). The $N_{32}$ contributions show up as dark purple bands in the FSB-$C_{\ell}$ blocks. Due to the weak signal in the generalised $gg\kappa$ and $g\kappa\kappa$ FSBs, this contribution may sometimes appear as a negative contribution (light blue pixels within the $N_{32}$ bands in the $\kappa-gggg$, gggggg-$g\kappa$ and $gg\kappa$-$g\kappa$ blocks).
@@ C.1.1 $gg\kappa\times gg\kappa$ covariance
We take a look at the “2+2+22+2+2” decomposition since we know from the single-field case (see Appendix C.1 of H25) that it should give rise to additional contributions along the main diagonal. To keep track of individual combinations of the 6 fields involved in this covariance, we write
@@
[…]
@@ C.2.1 $gg\kappa\times g\kappa$ covariance
for simplicity, and now look at the “3+23+2” decomposition which gives rise to noticeable off-diagonal contributions in the 1-field case. This gives us the 10 following combinations:
$\langle abcde\rangle_{32}=$ $\quad\;\cancel{\langle cde\rangle\langle ab\rangle}+\cancel{\langle bde\rangle\langle ac\rangle}+\langle bce\rangle\langle ad\rangle+\langle bcd\rangle\langle ae\rangle$
$+\cancel{\langle ade\rangle\langle bc\rangle}+\langle ace\rangle\langle bd\rangle+\langle acd\rangle\langle be\rangle$
@@ C.3 Calibrating the $f_{\mathrm{sky}}$ correction
This poses no issues in the 1-field case, where the original field and its filtered-squared version share the same footprint on the sky; the choice of rescaling factor becomes less obvious when the two fields each have their respective masks. Take Eq. 75 for instance; the expression includes $C^{gg}_{\ell}$ and $C^{g\kappa}_{\ell}$, which were measured on different effective sky areas – in this case, should the $N_{222}$ contribution be rescaled using $f_{\mathrm{sky}}=\langle w_{g}w_{g}\rangle$ or $f_{\mathrm{sky}}=\langle w_{g}w_{\kappa}\rangle$ ? The most conservative approach would be to always use the smallest effective sky fraction, as to not underestimate measurement errors, but this might be suboptimal.
@@
  • For the covariance of $\hat{\Phi}^{gg\kappa}_{LL\ell}$ and $\hat{\Phi}^{ggg}_{L^{\prime}L^{\prime}\ell^{\prime}}$, $N_{\tilde{\mathcal{S}}}$ is a subset of $N_{\mathcal{S}}$ by definition, since one power spectrum (gggggg FSB) is defined over the $g$ mask, while the other ($gg\kappa$) has for effective sky fraction the intersection of the $g$ and $\kappa$ masks. $N_{\tilde{\mathcal{S}}}$ and $N_{\mathcal{S}\cap\tilde{\mathcal{S}}}$ are therefore the same and cancel out, such that the expression must scale with $f_{\mathrm{sky}}=\langle w_{g}w_{g}\rangle$.
@@
Appendix D Interpolator Performance
We test the performance of the interpolator used in this paper, both in terms of accuracy and runtime. There are technically two stages in the bispectrum interpolator where we make use of interpolation:
  1. (i)
    First, instead of computing the bispectrum at each $(\ell_{1},\ell_{2},\ell_{3})$, we define a subset of multipoles, then interpolate linearly in 3D to evaluate missing configurations. The multipole subset we use here contains all multipoles between 0 and $10$, then every $20$ multipole from $10$ to $96$, and $50$ multipoles logarithmically spaced between $100$ and $\ell_{\mathrm{max}}$.
  2. (ii)
    We then build a cubic $\Omega_{m}$-interpolator, as the bispectrum $\Omega_{m}$ dependence is more complex and cannot be accounted for by a simple linear rescaling of the templates. We compute the FSB templates for $100$ values of $\Omega_{m}$ linearly spaced in the range $0.05\leq\Omega_{m}<0.5$; for a given value of $\Omega_{m}$, we can then interpolate the FSB from neighbouring values of $\Omega_{m}$.
The final interpolator is then a result of these two interpolations, at the level of the projected bispectrum templates (multipole interpolation), and at the FSB templates level ($\Omega_{m}$ interpolation).
We start by verifying that the overall accuracy of the interpolator is acceptable: we show in Fig. 13 that the interpolator works exquisitely well for the power spectra, and is slightly biased in the case of the bispectrum. Indeed, we report a difference between interpolation and full calculation of the FSB corresponding to $\sim 5\%$ of the errors on the gggggg measurement, which we deem acceptable. We investigate the source of this bias by computing the difference between the two approaches for each template contributing to the deterministic part of the gggggg FSB (see definitions in Eq. 27); we show this in the left panel of Fig. 14. The largest error on individual projected templates (at fixed $\Omega_{m}$ and $\sigma_{8}$, and before accounting for the bias rescaling) is found for the $T_{2}$ template with a bias of about $\lesssim 15\%$ of the full-calculation template. However, it is not the main contribution to the overall error budget, as the linear contribution $F_{2}$ accounts for most of the bispectrum signal. In blue, we also show the much smaller errors that could be achieved using a cubic interpolator at the multipole level (instead of a linear one, shown in red). In this case, the interpolation bias is barely noticeable, but the making of such cubic interpolators is very costly in 3D, even at small $\ell_{\text{max}}$. To generate $100$ FSB templates (i.e. 5 templates for each $\Omega_{m}$ value to build the interpolator), and allocating the same computational resources in both cases, the linear approach to interpolation takes about 4 minutes, but the cubic approach is closer to 38 hours. This could easily be sped up using parallelisation, and once built, the interpolator is fast to evaluate – whether it is linear or cubic; however, given our error budget and various approximations taken in this work, it is not yet necessary to refine the bispectrum interpolator to achieve sub-percent accuracy.
In the right panel of Fig. 14, we also show the time gain that can be obtained from using the interpolator. This gain evidently depends on the number of triangle contributions that need to be computed, which increases with the maximum multipole used in the analysis, $\ell_{\mathrm{max}}$. Since we selected a large-scale filter for the inference part of this work, the signal of the FSB dies off at reasonably low $\ell$, but the time gain is already consequential: to estimate a single $F_{2}$ template, for identical binning and cosmological parameters, the interpolator is 9 times faster to evaluate than the full bispectrum. However, we could have selected larger filters in the highest redshift bins – Fig. 4 makes this obvious – which would have then made us compute the bispectrum up to $\ell_{\mathrm{max}}\sim 300$. Reading from the plot, this corresponds to a $\sim\times 30$ improvement in runtime when using the interpolator over the full calculation [^10] We do not claim however that our implementation of the tree-level bispectrum calculation is fully optimized.. We do not show the same plot for $Q_{2}$ and $T_{2}$, since $F_{2}$ can partly be derived from these two templates – once you compute $F_{2}$, you effectively get $Q_{2}$ and $T_{2}$ for free if you save the various terms of $F_{2}$ separately (see Eq. 27). A full MCMC chain for the gggggg data vector with $50,000$ samples takes 0.6 hours to run on 8 cores, therefore we can expect a chain to take more than 5 hours to complete on the same number of cores using the full bispectrum calculation – which is not unfeasible given appropriate computing resources, but might become so if we push the calculation to smaller scales.
Figure 13: Interpolator accuracy. Left panel: We compare the power spectrum best-fit obtained from the interpolation (which comprises an interpolation at the multipole level, and an interpolation between values of $\Omega_{m}$) and the full calculation of the power spectrum. The difference between the two approaches is less than 0.003% of the error $\sigma$ associated with the measurement. Right panel: We show the same comparison for the gggggg FSB best-fit. The difference between the interpolation and full calculation corresponds in that case to about 5% of the error associated with the measurement. In both lower panels, we show in colour the interval corresponding to $10\%$ of the error bars in the upper panels.
Figure 14: Details by templates. Left panel: Interpolated best-fit template errors with respect to the full calculation, for the individual templates of the gggggg FSB. We do not show the interpolated noise templates, which depend on power spectra (which are themselves recovered to high precision by the interpolator, as shown in the left panel of Fig. 13). The red lines were obtained using linear interpolation between multipoles, while the blue lines show the equivalent for a cubic interpolator. Note how the right axis (cubic interpolation) is about $300$ times smaller than the left axis (linear interpolation). Right panel: Runtime performance of the interpolator (red dots) compared to computing the full calculation (black dots) for one $F^{ggg}_{LL\ell}$ template, as a function of maximum scale $\ell_{\mathrm{max}}$. As we increase $\ell_{\mathrm{max}}$, the number of triangle configurations contributing to the bispectrum dramatically increases, making the full calculation unsuitable for inference purposes. We highlight the data points relevant to this paper in a green box: for a scale cut of $\ell_{\mathrm{max}}=191$, interpolating is more than 99 times faster than computing the full bispectrum.